OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2.

OpenStudy (anonymous):

@Diyadiya @kyley_bear @happyrosy

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

f(5) = f(1)+f(2)+f(4) = 2+3+f(4) = 5+f(4) f(4) = f(1)+f(2) = f(3) = 5+f(3) f(3) = f(1) + f(2) + f(2) = 2+ 3 + 3=8 f(4) = 8+3 = 11 f(5) = 5+11=16

OpenStudy (zzr0ck3r):

make sense?

OpenStudy (anonymous):

Yeah Thank you!

OpenStudy (zzr0ck3r):

np